OpenStudy (anonymous):

-3(-x + 4) = -9 distribute then solve

OpenStudy (anonymous):

-3(-x+4)=-9 3x-12=-9 3x=-9+12 3x=3 x=1

OpenStudy (anonymous):

substitute the value of x to the equation -3(-1+4)=-9 -3(3)=-9 -9=-9

OpenStudy (anonymous):

insecure??why you should use this site just to say embarrasing word??